OpenStudy (anonymous):

I tryed doing this question but i keep doing it wrong. Solve |2 - 3x| = 7. A. {-5/3, 3} B. {-15, 3} C. {-3, 5/3} D. {-3, 15}

OpenStudy (lilg132):

it would help if you clicked good answer after people have answered it

OpenStudy (anonymous):

before you do anything, you have to get the numbers in the absolute value sign alone. In this equation, you do. So now you can just solve like a regular algebra problem. First problem will have what's NOT in the absolute value sign negative and positive. so first, you can do a problem where 7 is positive, and another problem where 7 is negative so 2-3x = -7 solve like a regular algebra problem -3x=-9 x=3 now make 7 positive 2-3x=7 solve regularly -3x=5 x= -5/3

OpenStudy (anonymous):

but it loooks like it's not on your list but im pretty sure it's right http://www.wolframalpha.com/input/?i=%7C2+-+3x%7C+%3D+7

OpenStudy (anonymous):

\[|3x-2|=7\] \[3x-2=7\text{ or } 3x-2=-7\] \[x=3 \text { or } x=-\frac{5}{3}\]